OpenStudy (anonymous):

Convert 13 + 7i to polar form

OpenStudy (anonymous):

z = √218 [ cos Θ + i sin Θ] where tan Θ = 7/13 ---> √218 e^(0.4939 i )

OpenStudy (anonymous):

Woah. Anyway you could clear up those ? marks?

OpenStudy (anonymous):

try refreshing the page

OpenStudy (anonymous):

Ah. That was strange. So it's... uh... √218(cos0.49 + isin0.49)?

OpenStudy (jdoe0001):

\(\large { \begin{array}{ccllll} 13 &+ {\color{blue}{ 7}}i\\ x&{\color{blue}{ y}} \end{array}\qquad \begin{cases} r=\sqrt{x^2+y^2}\\ \theta=tan^{-1}\left(\cfrac{y}{x}\right) \end{cases}\qquad r[cos(\theta)-isin(\theta)] }\)

OpenStudy (anonymous):

Oh wait, in degrees... 28.3

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

Thanks guys!

OpenStudy (jdoe0001):

yw